pooja195 (pooja195):

The admission fee at a small fair is $1.50 for children and $4.00 for adults. On a certain day, 2200 people enter the fair and $5050 is collected. How many children and how many adults attended? @e.mccormick @Nnesha This one im stuck on =(

Nnesha (nnesha):

x represent children and y represent adults children = x adults = y

sammixboo (sammixboo):

1.50x + 4y = 5050 and x + y = 2200

pooja195 (pooja195):

Thanks for the equations xD thats the part i needed help setting up

Nnesha (nnesha):

 all money tgther in one equation and for total number of children and adults x + y = total number of people